A wave with a frequency of 300 per second travels from water in which its velocity is 1450 m/s to air which its velocity is 331.5 m/s. Calculate the wavelengths in each medium


1
Expert's answer
2021-12-08T10:03:00-0500

"\lambda= \frac{v}{f}"

"\lambda_{watter}= \frac{1450}{300}\approx4,83m"


"\lambda_{air}= \frac{331.5}{300}\approx 1.11m"


"\text{Answer: }\lambda_{watter}= 4.83m; \lambda_{air}= 1.11m"